### v1

@@ 1 Introduction
Nucleus segmentation is an important task in computational pathology, with applications in cancer grading, tumor micro-environment analysis and survival prediction [van der Laak et al.(2021)van der Laak, Litjens, and Ciompi]. Most approaches rely on supervised learning, typically based on the UNet [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox] architecture. Several methods, such as StarDist [Weigert and Schmidt(2022)] and HoVerNet [Graham et al.(2019)Graham, Vu, Raza, Azam, Tsang, Kwak, and Rajpoot], have been developed to tackle domain-specific challenges, such as overlapping nuclei or joint segmentation and classification. Methods like HoVer-UNet [Tommasino et al.(2023)Tommasino, Russo, Rinaldi, and Ciompi], HoVerNeXt [Baumann et al.(2024)Baumann, Dislich, Rumberger, Nagtegaal, Martinez, and Zlobec], and InstanSeg [Goldsborough et al.(2024)Goldsborough, Philps, O’Callaghan, Inglis, Leplat, Filby, Bilen, and Bankhead] focus on computational efficiency without sacrificing segmentation accuracy. Despite these developments, existing models struggle to generalize to unseen conditions, for example due to different staining protocols, imaging devices, or sample extraction quality. Adapting them to new data involves training new models, which requires domain expertise to generate precise and exhaustive annotations (i.e. annotating each object in new patches for training). While tools like QuPath [Bankhead et al.(2017)Bankhead, Loughrey, Fernández, Dombrowski, McArt, Dunne, McQuaid, Gray, Murray, Coleman, James, Salto-Tellez, and Hamilton] are popular for this task, the procedure remains tedious and resource intensive.
@@
Recently, vision foundation models (VFM) for segmentation tasks have been introduced [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick, Zou et al.(2024)Zou, Yang, Zhang, Li, Li, Wang, Wang, Gao, and Lee]. They generalize to different imaging domains and support different segmentation tasks, including automatic and interactive segmentation. In particular, the Segment Anything Model (SAM) [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick] enables interactive segmentation from user input. SAM has been successfully applied and adapted to biomedical imaging domains [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape, Cheng et al.(2023)Cheng, Ye, Deng, Chen, Li, Wang, Su, Huang, Chen, Jiang, Sun, He, Zhang, Zhu, and Qiao, Yue et al.(2023)Yue, Zhang, Hu, Xia, Luo, and Wang, Archit et al.(2025)Archit, Freckmann, and Pape], but so far has not been comprehensively adapted to histopathology, where the performance of the original SAM is limited [Deng et al.(2023)Deng, Cui, Liu, Yao, Remedios, Bao, Landman, Wheless, Coburn, Wilson, Wang, Zhao, Fogo, Yang, Tang, and Huo, Zhang et al.(2023)Zhang, Ma, Kapse, Saltz, Vakalopoulou, Prasanna, and Samaras, Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape, Xu et al.(2024)Xu, Goetz, and Rajpoot]. Prior work used SAM for specific tasks in histopathology. For example, CellViT [Hörst et al.(2023)Hörst, Rempe, Heine, Seibold, Keyl, Baldini, Ugurel, Siveke, Grünwald, Egger, and Kleesiek] used its pretrained encoder for semantic instance segmentation. However, it only addresses automatic segmentation, neglecting interactive segmentation, and the available models are trained on a single dataset, PanNuke [Gamper et al.(2020)Gamper, Koohbanani, Benes, Graham, Jahanifar, Khurram, Azam, Hewitt, and Rajpoot]. Other approaches, such as [Xu et al.(2024)Xu, Goetz, and Rajpoot] use SAM for semi-supervised instance segmentation. However, this approach does not leverage the full potential of interactive segmentation with SAM, which enables human-in-the loop correction and data-specific finetuning [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape]. More recently, BioMedParse [Zhao et al.(2024)Zhao, Gu, Yang, Usuyama, Lee, Kiblawi, Naumann, Gao, Crabtree, Abel, Moung-Wen, Piening, Bifulco, Wei, Poon, and Wang], which is based on SEEM [Zou et al.(2024)Zou, Yang, Zhang, Li, Li, Wang, Wang, Gao, and Lee], has demonstrated text-prompt based semantic segmentation, but lacks support for instance segmentation. In summary, a VFM for automatic and interactive nucleus segmentation in histopathology is missing.
@@
Here, we present PathoSAM, a VFM for nucleus segmentation in histopathology. It supports interactive and automatic segmentation, based on finetuning SAM on a large and diverse annotated dataset. PathoSAM outperforms other SAM variants for interactive segmentation, is the new state-of-the-art for nucleus instance segmentation, and supports finetuning for semantic segmentation. PathoSAM is compatible with the user-friendly tools QuPath [Bankhead et al.(2017)Bankhead, Loughrey, Fernández, Dombrowski, McArt, Dunne, McQuaid, Gray, Murray, Coleman, James, Salto-Tellez, and Hamilton] and $\mu$SAM [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape]. We provide scripts to automatically segment whole-slide-images, together with the rest of our code, at https://github.com/computational-cell-analytics/patho-sam. Our contributions and a summary of results are shown in Fig. 1.
@@ 2.1 Segment Anything Model and Interactive Segmentation
SAM is trained on a large natural image dataset with annotations, using an objective that simulates interactive segmentation. In each iteration, this objective first samples prompts from an annotated mask, predicts the object mask, and then iteratively corrects the prediction with prompts sampled from the annotation. Predictions and annotations are compared with a loss function for each iteration, and the average loss is used for parameter updates. We use the training implementation of [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape].
@@ 2.2 Automatic Segmentation
$\mu$SAM [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape] introduces an alternative method for automatic instance segmentation, by adding a decoder, similar to UNETR [Hatamizadeh et al.(2021)Hatamizadeh, Tang, Nath, Yang, Myronenko, Landman, Roth, and Xu], to the model. It predicts foreground probabilities, the distance to the closest object center and the distance to the closest object boundary. During training, the decoder is jointly optimized with the rest of the model, using targets derived from ground-truth annotations. During inference, the decoder outputs are processed with a seeded watershed to obtain an instance segmentation. $\mu$SAM has termed this approach automatic instance segmentation (AIS), and has shown that it improves over AMG in terms of segmentation quality and computational efficiency for microscopy data. We evaluate interactive segmentation (for SAM or variants) with AMG and, if available, AIS, using the mean segmentation accuracy, (App. A).
@@ 2.3 A vision foundation model for histopathology
In order to train a VFM model for histopathology, we adopt the training procedure of $\mu$SAM for joint training of interactive (Sec.  2.1) and automatic (Sec.  2.2) segmentation. We train different models based on the datasets for nucleus segmentation (Sec.  2.5): three generalist models trained on 6 datasets, one for each of the three image encoder (ViT-B, ViT-L and ViT-H), which we call PathoSAM generalists, and a ViT-B model trained on PanNuke. The training settings can be found in App. C.
@@ 2.4 Semantic Segmentation
We extend PathoSAM to semantic segmentation, to support the analysis of different cell types. We add a separate decoder, following the same design as in Sec. 2.2. See also Fig. 4 for the combined architecture. This decoder predicts one output channel per semantic class, including background, and is optimized based on the cross entropy loss. We train and evaluate it on PanNuke ([Gamper et al.(2020)Gamper, Koohbanani, Benes, Graham, Jahanifar, Khurram, Azam, Hewitt, and Rajpoot], Sec. 2.5), which provides semantic annotations for 5 classes. Here, we compare 4 different training strategies, which start from a PathoSAM generalist: (i) freeze the image encoder, initialize the new decoder from scratch; (ii) finetune the image encoder initialize the new decoder from scratch; (iii) freeze the image encoder, initialize the new decoder with the AIS decoder weights; (iv) finetune the image encoder, initialize the new decoder with the AIS decoder weights. Semantic segmentation is evaluated with the class-frequency weighted dice score, see App. A.
@@ 2.5 Data
We assemble 14 publicly available histopathology datasets to train and evaluate PathoSAM. We use 6 datasets [Vu et al.(2019)Vu, Graham, Kurc, To, Shaban, Qaiser, Koohbanani, Khurram, Kalpathy-Cramer, Zhao, Gupta, Kwak, Rajpoot, Saltz, and Farahani, Graham et al.(2021)Graham, Jahanifar, Azam, Nimir, Tsang, Dodd, Hero, Sahota, Tank, Benes, Wahab, Minhas, Raza, Daly, Gopalakrishnan, Snead, and Rajpoot, Kumar et al.(2017)Kumar, Verma, Sharma, Bhargava, Vahadane, and Sethi, Gamper et al.(2020)Gamper, Koohbanani, Benes, Graham, Jahanifar, Khurram, Azam, Hewitt, and Rajpoot, Schuiveling et al.(2024)Schuiveling, Liu, Eek, Breimer, Suijkerbuijk, Blokx, and Veta, Naylor et al.(2019)Naylor, Laé, Reyal, and Walter] with tissue images stained with hematoxylin and eosin (H&E) staining and nucleus instance annotations to train our generalists. For each of these datasets we reserve a separate test split. An additional model is trained on PanNuke [Gamper et al.(2020)Gamper, Koohbanani, Benes, Graham, Jahanifar, Khurram, Azam, Hewitt, and Rajpoot], which is the most commonly used model for nucleus segmentation in histopathology. This dataset is also used for semantic segmentation, since it provides semantic annotations for 5 classes. We use the remaining 8 datasets [Graham et al.(2019)Graham, Vu, Raza, Azam, Tsang, Kwak, and Rajpoot, Mahbod et al.(2021)Mahbod, Schaefer, Bancher, Löw, Dorffner, Ecker, and Ellinger, Sirinukunwattana et al.(2016)Sirinukunwattana, Pluim, Chen, Qi, Heng, Guo, Wang, Matuszewski, Bruni, Sanchez, Böhm, Ronneberger, Cheikh, Racoceanu, Kainz, Pfeiffer, Urschler, Snead, and Rajpoot, Naji et al.(2024)Naji, Sancere, Simon, Büttner, Eich, Lohneis, and Bożek, Verma et al.(2021)Verma, Kumar, Patil, Kurian, Rane, Graham, Vu, Zwager, Raza, Rajpoot, Wu, Chen, Huang, Wang, Jung, Brown, Liu, Liu, Jahromi, Khani, Montahaei, Baghshah, Behroozi, Semkin, Rassadin, Dutande, Lodaya, Baid, Baheti, Talbar, Mahbod, Ecker, Ellinger, Luo, Dong, Xu, Yao, Lv, Feng, Xu, Zunair, Hamza, Smiley, Yin, Fang, Srivastava, Mahapatra, Trnavska, Zhang, Narayanan, Law, Yuan, Tejomay, Mitkari, Koka, Ramachandra, Kini, and Sethi, Alemi Koohbanani et al.(2020)Alemi Koohbanani, Jahanifar, Zamani Tajadin, and Rajpoot, Mahbod et al.(2024)Mahbod, Polak, Feldmann, Khan, Gelles, Dorffner, Woitek, Hatamikia, and Ellinger, Wang et al.(2024)Wang, Qiu, Hao, Jin, Gao, Qi, Xu, Zhang, and Xu] for out-of-domain evaluation, i.e to test our models on data which was not directly represented in the training set. Among these, two datasets contain different segmentation tasks, lymphocyte segmentation [Wang et al.(2024)Wang, Qiu, Hao, Jin, Gao, Qi, Xu, Zhang, and Xu] and gland segmentation [Sirinukunwattana et al.(2016)Sirinukunwattana, Pluim, Chen, Qi, Heng, Guo, Wang, Matuszewski, Bruni, Sanchez, Böhm, Ronneberger, Cheikh, Racoceanu, Kainz, Pfeiffer, Urschler, Snead, and Rajpoot]. They are used for separate evaluation experiments. Details on the datasets can be found in App. B.
@@
Figure 2: a) Automatic instance segmentation results for 8 datasets. Italic font means that the corresponding training split of the dataset was used for our generalist training, bold font means that it wasn’t. b) Interactive segmentation results for 8 datasets (same as in a)) for PathoSAM and other SAM variants. We report the interactive segmentation quality for initial point and box prompts as well as correction for 7 iterations after an initial point ($I_{P}$) and box ($I_{B}$).

### v2

@@ 1 Introduction
Nucleus segmentation is an important task in computational pathology, with applications in cancer grading, tumor micro-environment analysis and survival prediction [van der Laak et al.(2021)van der Laak, Litjens, and Ciompi]. Most approaches rely on supervised learning, typically based on UNets [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox]. Several methods, e.g. StarDist [Weigert and Schmidt(2022)] or HoVerNet [Graham et al.(2019)Graham, Vu, Raza, Azam, Tsang, Kwak, and Rajpoot], address domain-specific challenges, such as overlapping nuclei or joint segmentation and classification. HoVer-UNet [Tommasino et al.(2024)Tommasino, Russo, Rinaldi, and Ciompi], HoVerNeXt [Baumann et al.(2024)Baumann, Dislich, Rumberger, Nagtegaal, Martinez, and Zlobec], and InstanSeg [Goldsborough et al.(2024)Goldsborough, Philps, O’Callaghan, Inglis, Leplat, Filby, Bilen, and Bankhead] focus on efficiency without sacrificing segmentation accuracy. Despite these developments, existing models struggle to generalize to new conditions, for example due to different staining protocols, imaging devices, or sample extraction quality. Adapting them to new data involves training new models, which requires domain expertise for precise and exhaustive annotations (drawing each nucleus in new training patches). While tools like QuPath [Bankhead et al.(2017)Bankhead, Loughrey, Fernández, Dombrowski, McArt, Dunne, McQuaid, Gray, Murray, Coleman, James, Salto-Tellez, and Hamilton] are popular for this task, the procedure remains tedious and resource intensive.
@@
Recently, vision foundation models (VFM) for segmentation tasks have been introduced [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick, Zou et al.(2024)Zou, Yang, Zhang, Li, Li, Wang, Wang, Gao, and Lee]. They generalize to different imaging domains and support different segmentation tasks, including automatic and interactive segmentation. In particular, the Segment Anything Model (SAM) [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick] enables interactive segmentation from user input. SAM has been successfully applied and adapted to biomedical imaging domains [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape, Cheng et al.(2023)Cheng, Ye, Deng, Chen, Li, Wang, Su, Huang, Chen, Jiang, Sun, He, Zhang, Zhu, and Qiao, Yue et al.(2024)Yue, Zhang, Hu, Xia, Luo, and Wang, Archit et al.(2025b)Archit, Freckmann, and Pape], but so far has not been comprehensively adapted to histopathology, where the performance of the original SAM is limited [Deng et al.(2023)Deng, Cui, Liu, Yao, Remedios, Bao, Landman, Tang, Wheless, Coburn, Wilson, Wang, Fogo, Yang, and Huo, Zhang et al.(2023)Zhang, Ma, Kapse, Saltz, Vakalopoulou, Prasanna, and Samaras, Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape, Xu et al.(2024)Xu, Goetz, and Rajpoot]. Prior work used SAM for specific tasks in histopathology. For example, CellViT [Hörst et al.(2024)Hörst, Rempe, Heine, Seibold, Keyl, Baldini, Ugurel, Siveke, Grünwald, Egger, and Kleesiek] used its pretrained encoder for semantic instance segmentation. However, it only addresses automatic segmentation, neglecting interactive segmentation, and the available models are trained on a single dataset, PanNuke [Gamper et al.(2019)Gamper, Koohbanani, Benes, Khuram, and Rajpoot]. Other approaches, such as [Xu et al.(2024)Xu, Goetz, and Rajpoot] use SAM for semi-supervised instance segmentation. However, this approach does not leverage the full potential of interactive segmentation with SAM, which enables human-in-the-loop correction and data-specific finetuning [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape]. More recently, BioMedParse [Zhao et al.(2024)Zhao, Gu, Yang, Usuyama, Lee, Kiblawi, Naumann, Gao, Crabtree, Abel, Moung-Wen, Piening, Bifulco, Wei, Poon, and Wang], which is based on SEEM [Zou et al.(2024)Zou, Yang, Zhang, Li, Li, Wang, Wang, Gao, and Lee], has demonstrated text-prompt based semantic segmentation, but lacks support for instance segmentation. In summary, a VFM for automatic and interactive nucleus segmentation in histopathology is missing.
@@
Here, we present PathoSAM, a VFM for nucleus segmentation in histopathology. It supports interactive and automatic segmentation, based on finetuning SAM on a large and diverse annotated dataset. PathoSAM outperforms other SAM variants for interactive segmentation, is the new state-of-the-art for nucleus instance segmentation, and supports finetuning for semantic segmentation. PathoSAM is compatible with the user-friendly tools QuPath [Bankhead et al.(2017)Bankhead, Loughrey, Fernández, Dombrowski, McArt, Dunne, McQuaid, Gray, Murray, Coleman, James, Salto-Tellez, and Hamilton] and $\mu$SAM [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape]. We provide scripts to automatically segment whole-slide-images, together with the rest of our code, at https://github.com/computational-cell-analytics/patho-sam. Our contributions and a summary of results are shown in Fig. 1.
@@ 2.1 Segment Anything Model and Interactive Segmentation
SAM is trained on a large natural image dataset with annotations, using an objective that simulates interactive segmentation. In each iteration, this objective first samples prompts from an annotated mask, predicts the object mask, and then iteratively corrects the prediction with prompts sampled from the annotation. Predictions and annotations are compared with a loss function for each iteration, and the average loss is used for parameter updates. We use the training implementation of $\mu$SAM [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape].
@@ 2.2 Automatic Segmentation
$\mu$SAM [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape] introduces an alternative method for automatic instance segmentation, by adding a decoder, similar to UNETR [Hatamizadeh et al.(2022)Hatamizadeh, Tang, Nath, Yang, Myronenko, Landman, Roth, and Xu], to the model. It predicts foreground probabilities, the distance to the closest object center and the distance to the closest object boundary. During training, the decoder is jointly optimized with the rest of the model, using targets derived from ground-truth annotations. During inference, the decoder outputs are processed with a seeded watershed to obtain an instance segmentation. $\mu$SAM has termed this approach automatic instance segmentation (AIS), and has shown that it improves over AMG in terms of segmentation quality and computational efficiency for microscopy data. We evaluate interactive segmentation (for SAM or variants) with AMG and, if available, AIS, using the mean segmentation accuracy, (App. A).
@@ 2.3 A vision foundation model for histopathology
In order to train a VFM model for histopathology, we adopt the training procedure of $\mu$SAM for joint training of interactive (Sec.  2.1) and automatic (Sec.  2.2) segmentation. We train different models based on the datasets for nucleus segmentation (Sec.  2.5): three generalist models trained on 6 datasets, one for each of the three image encoders (ViT-B, ViT-L and ViT-H), which we call PathoSAM generalists, and a ViT-B model trained on PanNuke. The training settings can be found in App. C.
@@ 2.4 Semantic Segmentation
We extend PathoSAM to semantic segmentation, to support the analysis of different cell types. We add a separate decoder, following the same design as in Sec. 2.2. See also Fig. 5 for the combined architecture. This decoder predicts one output channel per semantic class, including background, and is optimized based on the cross entropy loss. We train and evaluate it on PanNuke ([Gamper et al.(2019)Gamper, Koohbanani, Benes, Khuram, and Rajpoot], Sec. 2.5), which provides semantic annotations for 5 classes. Here, we compare 4 different training strategies, which start from a PathoSAM generalist: (i) freeze the image encoder, initialize the new decoder from scratch; (ii) finetune the image encoder initialize the new decoder from scratch; (iii) freeze the image encoder, initialize the new decoder with the AIS decoder weights; (iv) finetune the image encoder, initialize the new decoder with the AIS decoder weights. Semantic segmentation is evaluated with the class-frequency weighted dice score, see App. A.
@@ 2.5 Data
We assemble 14 publicly available histopathology datasets to train and evaluate PathoSAM. We use 6 datasets [Vu et al.(2019)Vu, Graham, Kurc, To, Shaban, Qaiser, Koohbanani, Khurram, Kalpathy-Cramer, Zhao, Gupta, Kwak, Rajpoot, Saltz, and Farahani, Graham et al.(2021)Graham, Jahanifar, Azam, Nimir, Tsang, Dodd, Hero, Sahota, Tank, Benes, Wahab, Minhas, Raza, El Daly, Gopalakrishnan, Snead, and Rajpoot, Kumar et al.(2017)Kumar, Verma, Sharma, Bhargava, Vahadane, and Sethi, Gamper et al.(2019)Gamper, Koohbanani, Benes, Khuram, and Rajpoot, Schuiveling et al.(2025)Schuiveling, Liu, Eek, Breimer, Suijkerbuijk, Blokx, and Veta, Naylor et al.(2019)Naylor, Laé, Reyal, and Walter] with tissue images stained with hematoxylin and eosin (H&E) staining and nucleus instance annotations to train our generalists. For each of these datasets we reserve a separate test split. An additional model is trained on PanNuke [Gamper et al.(2019)Gamper, Koohbanani, Benes, Khuram, and Rajpoot], which is the most popular dataset for nucleus segmentation. This dataset is also used for semantic segmentation, since it provides semantic annotations for 5 classes. We use the remaining 8 datasets [Graham et al.(2019)Graham, Vu, Raza, Azam, Tsang, Kwak, and Rajpoot, Mahbod et al.(2021)Mahbod, Schaefer, Bancher, Löw, Dorffner, Ecker, and Ellinger, Sirinukunwattana et al.(2017)Sirinukunwattana, Pluim, Chen, Qi, Heng, Guo, Wang, Matuszewski, Bruni, Sanchez, Böhm, Ronneberger, Cheikh, Racoceanu, Kainz, Pfeiffer, Urschler, Snead, and Rajpoot, Naji et al.(2024)Naji, Sancere, Simon, Büttner, Eich, Lohneis, and Bożek, Verma et al.(2021)Verma, Kumar, Patil, Kurian, Rane, Graham, Vu, Zwager, Raza, Rajpoot, Wu, Chen, Huang, Wang, Jung, Brown, Liu, Liu, Jahromi, Khani, Montahaei, Baghshah, Behroozi, Semkin, Rassadin, Dutande, Lodaya, Baid, Baheti, Talbar, Mahbod, Ecker, Ellinger, Luo, Dong, Xu, Yao, Lv, Feng, Xu, Zunair, Hamza, Smiley, Yin, Fang, Srivastava, Mahapatra, Trnavska, Zhang, Narayanan, Law, Yuan, Tejomay, Mitkari, Koka, Ramachandra, Kini, and Sethi, Alemi Koohbanani et al.(2020)Alemi Koohbanani, Jahanifar, Zamani Tajadin, and Rajpoot, Mahbod et al.(2024)Mahbod, Polak, Feldmann, Khan, Gelles, Dorffner, Woitek, Hatamikia, and Ellinger, Wang et al.(2024)Wang, Qiu, Hao, Jin, Gao, Qi, Xu, Zhang, and Xu] for out-of-domain evaluation, i.e to test our models on data which was not directly represented in the training set. Among these, two datasets contain different segmentation tasks, lymphocyte segmentation [Wang et al.(2024)Wang, Qiu, Hao, Jin, Gao, Qi, Xu, Zhang, and Xu] and gland segmentation [Sirinukunwattana et al.(2017)Sirinukunwattana, Pluim, Chen, Qi, Heng, Guo, Wang, Matuszewski, Bruni, Sanchez, Böhm, Ronneberger, Cheikh, Racoceanu, Kainz, Pfeiffer, Urschler, Snead, and Rajpoot]. They are used for separate evaluation experiments. Details on the datasets can be found in App. B.
@@
Figure 2: a) Automatic instance segmentation results for 8 datasets. Italic font indicates that the corresponding training split of the dataset was used for our generalist training, while bold font indicates that it was not. b) Interactive segmentation results for 8 datasets (same as in a)) for PathoSAM and other SAM variants. We report the interactive segmentation quality for initial point and box prompts as well as correction for 7 iterations after an initial point ($I_{P}$) and box ($I_{B}$).